Superman leaps in front of Lois Lane to save her from a volley of bullets. In a 1-minute interval, an automatic weapon fires 156bullets, each of mass 8.0 g, at 360 m/s. The bullets strike his mighty chest, which has an area of 0.71 m2. Find the average force exerted on Superman's chest if the bullets bounce back after an elastic, head-on collision.
N
Answer #1

mass of the "all bullets" is =156*8 =1248 grams=1.248 kg.

Their total change in velocity is 720m/s (because its 360 - (-360)) .

Ergo, acceleration equals 720/60 sec = 12 m/s^2.

F = ma,

force = 1.248 * 12 = 14.976 Newtons.
Answer #2

the mass of the "bullets" (we will treat them as a single object) is 1.248 kg. Their total change in velocity is 720 m/s (because its 360 - (-360.)) Ergo, acceleration equals 720/60 sec = 12 m/s^2. F = ma, so force = 1.248* 12 = 14.976 Newtons.